The number of ballots cast a parish election is 12,969.if the number of registered voters in the parish is 19,800, what percent has voted
Found 2 solutions by stanbon, ewatrrr: Answer by stanbon(75887) 
 The number of ballots cast a parish election is 12,969.if the number of registered voters in the parish is 19,800, what percent has voted
---------------
x/100 = 12,969/19,800
--------
x/100 = 0.655
---
x = 65.6%
==============
Cheers,
Stan H.
==============
Answer by ewatrrr(24785) 
| |
|